the no. of side two regular polygon are as5:4 and the difference between their angles is 9 degree .find the no. of sides of the polygon

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

for n sides, each angle is (1-(2/n))180 degrees =x degrees,for 5n/4 sides, each angle is (1-(2/(5n/4)))180 degrees=y degrees, y-x=9 degrees, solving the resulting elementary equation in 'n' we get number of sides = 8 and 10